Suppose a single firm produces all of the output in a contestable market. The market inverse demand function is P = 200 -5Q, and the firm’s cost function is C(Q) = 20Q. Determine the firm’s equilibrium price and corresponding profits.
Answer #1

From the given information the marginal revenue function is MR = 200 - 10Q and marginal cost function is MC = 20

Use MR = MC

200 - 10Q = 20

Q = 180/10 = 18 units

P = 200-5*18 =110

Profit = revenue - cost = 18 x 110 - 18 x 20 = $1620

Therefore the equilibrium price is $110 per unit and the corresponding profit is $1620